# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v2

@@ Acknowledgements.
We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback and contributions.References
@@
- AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
- AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
@@
- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. https://www.anthropic.com/news/visible-extended-thinking
@@
- Chen et al. (2023) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W. Cohen. 2023. Program of Thoughts Prompting: Disentangling Computation from Reasoning for Numerical Reasoning Tasks. Transactions on Machine Learning Research (2023).
@@
- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
- guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
@@
- Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
@@
- Lyu et al. (2023) Qing Lyu, Shreya Havaldar, Adam Stein, Li Zhang, Delip Rao, Eric Wong, Marianna Apidianaki, and Chris Callison-Burch. 2023. Faithful Chain-of-Thought Reasoning. arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
@@
- Malkov and Yashunin (2020) Yu A. Malkov and D. A. Yashunin. 2020. Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs. IEEE Transactions on Pattern Analysis and Machine Intelligence 42, 4 (2020), 824–836. https://doi.org/10.1109/TPAMI.2018.2889473
@@
- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. https://openai.com/index/introducing-text-and-code-embeddings/
- OpenAI (2024) OpenAI. 2024. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
@@
- TruLens (2021) TruLens. 2021. TruLens. https://www.trulens.org/

### v3

@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant Nos 2439559 and 2040433. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback.References
@@
- Agarwal and Huang (2025) Arun Agarwal and Renee Huang. 2025. Introducing Cortex AISQL: Reimagining SQL into AI Query Language for Multimodal Data. Retrieved 2025-07-08 from https://www.snowflake.com/en/blog/ai-sql-query-language/
- AI (2023a) Guardrails AI. 2023a. Guardrails. Retrieved 2025-07-08 from https://github.com/guardrails-ai/guardrails
- AI (2023b) LangChain AI. 2023b. LangChain. Retrieved 2025-07-08 from https://github.com/langchain-ai/langchain
@@
- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. Retrieved 2025-07-08 from https://www.anthropic.com/news/visible-extended-thinking
@@
- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. Retrieved 2025-07-08 from https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
- guidance (2022) guidance. 2022. guidance-ai. Retrieved 2025-07-08 from https://github.com/guidance-ai/guidance
@@
- Index (2023) Llama Index. 2023. Llama Index. Retrieved 2025-07-08 from https://github.com/run-llama/llama_index
@@
- Loula et al. (2025) João Loula, Benjamin LeBrun, Li Du, Ben Lipkin, Clemente Pasti, Gabriel Grand, Tianyu Liu, Yahya Emara, Marjorie Freedman, Jason Eisner, Ryan Cotterell, Vikash Mansinghka, Alexander K. Lew, Tim Vieira, and Timothy J. O’Donnell. 2025. Syntactic and Semantic Control of Large Language Models via Sequential Monte Carlo. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=xoXn62FzD0
@@
- Madden et al. (2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska. 2024. Databases Unbound: Querying All of the World’s Bytes with AI. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4546–4554. https://doi.org/10.14778/3685800.3685916
@@
- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. Retrieved 2025-07-08 from https://openai.com/index/introducing-text-and-code-embeddings/
- OpenAI (2024) OpenAI. 2024. Structured Outputs. Retrieved 2025-07-08 from https://platform.openai.com/docs/guides/structured-outputs/
- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. Retrieved 2025-07-08 from https://github.com/json-schema-org/json-schema-spec
@@
- TruLens (2021) TruLens. 2021. TruLens. Retrieved 2025-07-08 from https://www.trulens.org/